# Interpretable Oracle Bone Script Decipherment through Radical and Pictographic Analysis with LVLMs

### v1

@@ Introduction
Our framework is built upon Qwen2.5-VL (Bai et al. 2025), sharing the same vision encoder and LLM. As illustrated in Figure 3, we introduce a spatial patch merger as the visual adapter and a classifier to predict the radical label. We also propose a radical LoRA module $\phi_{\mathrm{rad}}$ and a pictographic LoRA module $\phi_{\mathrm{pic}}$ (Hu et al. 2021) to analyze the corresponding information. Furthermore, we design a progressive training—starting from radical recognition, followed by radical and pictographic analysis, and culminating in mutual analysis—to gradually lead the model to OBS decipherment task. In addition, we propose a novel radical-pictographic dual matching mechanism to select the most appropriate character from the database.
@@
First, we filter candidates by the predicted radical label $\hat{\mathbf{r}}$, then select the top-$k$ entries by the semantic similarity $\mathcal{S}(\mathbf{a}_{\mathrm{pic},i},\hat{\mathbf{a}}_{\mathrm{pic}})$ (BERT-Score) (Devlin et al. 2019) between the pictographic analyses. Second, we concatenate the predicted radical and radical-informed pictographic analyses, and select another top-$k$ entries by similarity $\mathcal{S}((\mathbf{a}_{\mathrm{rad},i}\oplus\mathbf{a}_{\mathrm{joint},i}),(\hat{\mathbf{a}}_{\mathrm{rad}}\oplus\hat{\mathbf{a}}_{\mathrm{joint}}))$. Finally, we merge and rerank these candidate sets to obtain the top-$k$ modern Chinese characters as decipherment results. All steps and notations are detailed in Algorithm 1.
@@
To quantitatively evaluate the precision of radical and pictographic analysis produced by our method, we employ BERT-Score (Devlin et al. 2019) to measure the similarity between the top-1 outputs and the ground-truth labels from the dictionary $\mathcal{D}$. We also evaluate other LVLMs, including GPT-4.1, Qwen-VL-Max, and Qwen2.5-VL-7B, and compare their average BERT-Score on both the validation and zero-shot settings of the HUST-OBC (Wang et al. 2024a) and EVOBC (Wang et al. 2024c) datasets. As shown in Table 2, our method significantly outperforms the state-of-the-art LVLM GPT-4.1 21.60% and 12.95%, averaged over the two datasets, under validation and zero-shot settings, respectively. This result indicates that the analyses generated by our framework are more reliable and informative.
@@
To evaluate the effectiveness of the proposed radical recognition stage, we use the original vision encoder of Qwen2.5-VL-7B (Bai et al. 2025) as the baseline, and incrementally incorporate our radical recognition module as well as a LoRA-based recognition method. Their recognition accuracy is reported on the HUST-OBS dataset (Wang et al. 2024a) with validation and zero-shot settings. Our method introduces a spatial patch merger and the loss function $\mathcal{L}_{\mathrm{trip}}$ on top of the baseline vision encoder, resulting in improvements of 0.9% and 1.2% accuracy on the validation and zero-shot settings, respectively. The LoRA-based recognition method merges the recognition stage with the radical analysis process and training with LoRA-based fine-tuning. The results demonstrate that this method leads to a significant drop in radical recognition accuracy and introduces substantial errors in radical analysis, thus we retain radical recognition as an independent stage in our framework.
@@
Figure 5 presents qualitative results of our method and OBSD (Guan et al. 2024) under three settings: validation, zero-shot, and undeciphered OBS. As shown, our model demonstrates strong recognition capabilities on the validation set and also generalizes well to unseen OBS in zero-shot settings. More notably, for characters that remain undeciphered by human experts, the model is able to produce semantically plausible predictions, accompanied by interpretable analyses. Our design of Radical-Pictographic Mutual Analysis plays a key role: The radical analysis component traces the structural origins of radicals and explains their symbolic function in the current character form. Meanwhile, the pictographic form analysis provides a holistic visual-semantic mapping based on the character’s overall shape and implied meaning. Together, these complementary analyses form a dual reasoning path that enhances the model’s ability to generate semantically grounded and interpretable outputs, even for previously untranscribed scripts.
@@ Conclusion
We propose an interpretable OBS decipherment framework through radical and pictographic analysis. The framework bridges glyph to meaning through three stages: radical recognition and analysis, pictographic analysis, and mutual analysis. With the proposed radical-pictographic dual matching, our model can filter the appropriate deciphering candidate set from the dictionary based on analysis results, replacing direct output of deciphering results to achieve better zero-shot performance. Moreover, the generated textual analyses serve as interpretable content, offering references for undeciphered OBS characters, thus holding great potential for archaeological applications. To support training, we construct the PD-OBS dataset, comprising 47,157 Chinese characters annotated with OBS images and pictographic analysis texts, providing a valuable resource for future research. Experimental results demonstrate that our method achieves strong performance in decipherment accuracy, generalization, and interoperability.

### v2

@@ Introduction
Our framework is built upon Qwen2.5-VL-7B (Bai et al. 2025), sharing the same vision encoder and LLM. As illustrated in Figure 3, we introduce a spatial patch merger as the visual adapter and a classifier to predict the radical label. We also propose a radical LoRA module $\phi_{\mathrm{rad}}$ and a pictographic LoRA module $\phi_{\mathrm{pic}}$ (Hu et al. 2021) to analyze the corresponding information. Furthermore, we design a progressive training—starting from radical recognition, followed by radical and pictographic analysis, and culminating in mutual analysis—to gradually lead the model to OBS decipherment task. In addition, we propose a novel radical-pictographic dual matching mechanism to select the most appropriate character from the database.
@@
First, we filter candidates by the predicted radical label $\hat{\mathbf{r}}$, then select the top-$k$ entries by the semantic similarity $\mathcal{S}(\mathbf{a}_{\mathrm{pic},i},\hat{\mathbf{a}}_{\mathrm{pic}})$ (BERT-Score) (Devlin et al. 2019) between the pictographic analysis. Second, we concatenate the radical analysis and radical-informed pictographic analysis, and select another top-$k$ entries by similarity $\mathcal{S}((\mathbf{a}_{\mathrm{rad},i}\oplus\mathbf{a}_{\mathrm{joint},i}),(\hat{\mathbf{a}}_{\mathrm{rad}}\oplus\hat{\mathbf{a}}_{\mathrm{joint}}))$. Finally, we merge and rerank these candidate sets to obtain the top-$k$ modern Chinese characters as decipherment results. All steps and notations are detailed in Algorithm 1.
@@
To quantitatively evaluate the precision of radical and pictographic analysis produced by our method, we employ BERT-Score (Devlin et al. 2019) to measure the similarity between the top-1 outputs and the ground-truth labels from the dictionary $\mathcal{D}$. We also evaluate other LVLMs, including GPT-4.1, Qwen-VL-Max, and Qwen2.5-VL-7B, and compare their average BERT-Score on both the validation and zero-shot settings of the HUST-OBC (Wang et al. 2024a) and EVOBC (Wang et al. 2024c) datasets. As shown in Table 2, our method significantly outperforms the state-of-the-art LVLM GPT-4.1 21.60% and 12.95%, averaged over the two datasets, under validation and zero-shot settings, respectively. This result indicates that the analysis generated by our framework are more reliable and informative.
@@
To evaluate the effectiveness of the proposed radical recognition stage, we use the original vision encoder of Qwen2.5-VL-7B (Bai et al. 2025) as the baseline, and incorporate our radical recognition module or a LoRA-based recognition method. Their recognition accuracy is reported on the HUST-OBS dataset (Wang et al. 2024a) with validation and zero-shot settings. Our method introduces a spatial patch merger and the loss function $\mathcal{L}_{\mathrm{trip}}$ on top of the baseline vision encoder, resulting in improvements of 0.9% and 1.2% accuracy on the validation and zero-shot settings, respectively. The LoRA-based recognition method merges the recognition stage with the radical analysis process and training with LoRA-based fine-tuning. The results demonstrate that this method leads to a significant drop in radical recognition accuracy and introduces substantial errors in radical analysis, thus we retain radical recognition as an independent stage in our framework.
@@
Figure 5 presents qualitative results of our method and OBSD (Guan et al. 2024) under three settings: validation, zero-shot, and undeciphered OBS. As shown, our model demonstrates strong recognition capabilities on the validation set and also generalizes well to unseen OBS in zero-shot settings. More notably, for characters that remain undeciphered by human experts, the model is able to produce semantically plausible predictions, accompanied by interpretable analysis. Our design of Radical-Pictographic Mutual Analysis plays a key role: The radical analysis component traces the structural origins of radicals and explains their symbolic function in the current character form. Meanwhile, the pictographic form analysis provides a holistic visual-semantic mapping based on the character’s overall shape and implied meaning. Together, these complementary analysis form a dual reasoning path that enhances the model’s ability to generate semantically grounded and interpretable outputs, even for previously untranscribed scripts.
@@ Conclusion
We propose an interpretable OBS decipherment framework through radical and pictographic analysis. The framework bridges glyph to meaning through three stages: radical recognition and analysis, pictographic analysis, and mutual analysis. With the proposed radical-pictographic dual matching, our model can filter the appropriate deciphering candidate set from the dictionary based on analysis results, replacing direct output of deciphering results to achieve better zero-shot performance. Moreover, the generated textual analysis serve as interpretable content, offering references for undeciphered OBS characters, thus holding great potential for archaeological applications. To support training, we construct the PD-OBS dataset, comprising 47,157 Chinese characters annotated with OBS images and pictographic analysis texts, providing a valuable resource for future research. Experimental results demonstrate that our method achieves strong performance in decipherment accuracy, generalization, and interpretability.